OpenStudy (anonymous):

help me pls. it's our homework in integral calculus. find the indefinite integral of the ff. using inverse trigonometric functions: 1. integral of {1/[(sqrtx)(sqrt(1-x))]} dx 2. integral of [2x/(x^2 + 6x + 3)] dx 3. integral of [x+5 /sqrt (9-(x-3)^2)] dx

ganeshie8 (ganeshie8):

for #1, complete the square in denominator

OpenStudy (anonymous):

here are the "clear" eqns: \[1. \int\limits_{}^{} \frac{ 1 }{ \sqrt{x} \sqrt{1-x} } dx\] \[2. \int\limits_{}^{} \frac{ 2x }{ x^{2} + 6x + 3 } dx\] \[3. \int\limits_{}^{} \frac{ x + 5 }{ \sqrt{9 - (x - 3)^{2}} } dx\]

OpenStudy (anonymous):

@ganeshie8

random231 (random231):

 complete the square